Pinnacle Plus declared and paid a cash dividend of $9,100 in the current year. Its comparative financial statements, prepared at December 31, reported the following summarized information:

Current Year Previous Year
Income Statement
Sales Revenue $ 235,000 $ 199,000
Cost of Goods Sold 102,000 98,000
Gross Profit 133,000 101,000
Operating Expenses 61,000 53,000
Interest Expense 6,500 6,500
Income before Income Tax Expense 65,500 41,500
Income Tax Expense (30%) 19,650 12,450
Net Income $ 45,850 $ 29,050
Balance Sheet
Cash $ 102,125 $ 13,000
Accounts Receivable, Net 42,000 37,000
Inventory 50,000 63,000
Property and Equipment, Net 120,000 130,000
Total Assets $ 314,125 $ 243,000
Accounts Payable $ 67,000 $ 32,500
Income Tax Payable 1,625 1,750
Note Payable (long-term) 65,000 65,000
Total Liabilities 133,625 99,250
Common Stock (par $10) 105,000 105,000
Retained Earnings 75,500 38,750
Total Liabilities and Stockholders’ Equity $ 314,125 $ 243,000

Required:

  1. Compute the gross profit percentage in the current and previous years. Are the current year results better, or worse, than those for the previous year?
  2. Compute the net profit margin for the current and previous years. Are the current year results better, or worse, than those for the previous year?
  3. Compute the earnings per share for the current and previous years. Are the current year results better, or worse, than those for the previous year?
  4. Stockholders’ equity totaled $125,000 at the beginning of the previous year. Compute the return on equity (ROE) ratios for the current and previous years. Are the current year results better, or worse, than those for the previous year?
  5. Net property and equipment totaled $135,000 at the beginning of the previous year. Compute the fixed asset turnover ratios for the current and previous years. Are the current year results better, or worse, than those for the previous year?
  6. Compute the debt-to-assets ratios for the current and previous years. Is debt providing financing for a larger or smaller proportion of the company’s asset growth?
  7. Compute the times interest earned ratios for the current and previous years. Are the current year results better, or worse, than those for the previous year?
  8. After Pinnacle Plus released its current year’s financial statements, the company’s stock was trading at $43. After the release of its previous year’s financial statements, the company’s stock price was $40 per share. Compute the P/E ratios for both years. Does it appear that investors have become more (or less) optimistic about Pinnacle’s future success?
Answer #1

Answer:

1.Gross profit percentage = Gross Profit / Sales

Current Year = 133000 / 235000 = 56.60% ; Previous Year = 101000/199000 = 50.75%

Current Year results are better than previous year.

2.Net Profit Margin = Net Income / Sales

Current Year = 45850 / 235000 = 19.51% ; Previous Year = 29050 /199000 = 14.60%

Current Year results are better than previous year.

3.Earning per share = Net Income / Outstanding shares

Current Year = 45850 / 10500 = 4.37 ; Previous Year = 29050 / 10500 = 2.77

Current Year results are better than previous year.

4.ROE = Net Income / Avg. Stockholders Equity

Current Year = 45850 / (105000 + 105000 + 75500 + 38750 ) / 2 = 45850 / 162125 = 0.28

Previous Year = 29050 / (105000 + 38750 + 125000 ) / 2 = 29050 / 134375 = 0.216

Current Year results are better than previous year.

5.Fixed Asset Turnover ratio = Sales Revenue / Avg.Fixed Assets

Current Year = 235000 / (120000 + 130000 ) / 2 = 235000 / 125000 = 1.88

Previous Year = 199000 / (130000 + 135000 ) / 2 = 199000 / 132500 = 1.50

Current Year results are better than previous year.

6.Debt to Assets Ratio = Total Liabilities / Total Assets

Current Year = 133625 / 314125 = 0.425

Previous Year = 99250 / 243000 = 0.408

Current Year results are better than previous year.

7.Times Interest earned = Operating Income / Interest Expense

Current Year = (133000 - 61000 ) / 6500 = 11.08

Previous Year = (101000 - 53000 ) / 6500 = 7.38

Current Year results are better than previous year.

8.P/E Ratio = Market Price / EPS

Current Year = 43 / 4.37 = 9.84 ; Previous Year = 40 / 2.77 = 14.44